OpenStudy (anonymous):

Which of the following fractions are equivalent to 6/10 1. 3/5 2.12/18 3.2/3 4. 9/15

OpenStudy (anonymous):

6/2 = 3 10/2 = 5 Therefore 3/5 = 6/10!